OpenStudy (anonymous):

Choose the correct x-intercept and y-intercept of the equation x + 2y = 4. x-intercept: (0, 2), y-intercept: (4, 0) x-intercept: (2, 0), y-intercept: (4, 0) x-intercept: (0, 4), y-intercept: (2, 0) x-intercept: (4, 0), y-intercept: (0, 2)

OpenStudy (anonymous):

x-intercept occurs at y=0, y-intercept occurs at x=0

OpenStudy (anonymous):

4th one

OpenStudy (anonymous):

d

OpenStudy (anonymous):

thanks

OpenStudy (ash2326):

standar equation of a line with x intercept (a,0) and y-intercept (0, b) \[\frac{x}{a}+\frac{y}{b}=1\] here the equation is \[x+2y=4\]divide the whole equation by 4 \[\frac{x}{4}+\frac{y}{2}=1\] so xintercept (4, 0) and y intercept (0, 2). Option D

OpenStudy (anonymous):

i believe it is (D)

OpenStudy (anonymous):

thanks for explaning @ash